OpenStudy (anonymous):

How do you determine the domain and sketch the graph of the reducible function for: f(x) = (x + 2) / (x^2 -x - 6 )

OpenStudy (anonymous):

The domain of the original function is x is all reals except -2 and 3. If you factor the denominator you get f(x) = (x + 2) / ((x-3)(x+2)) cancel out the x+2 terms f(x) = 1/(x-3) If you graph this function, your original function is the same except it has a hole at x = -2.

OpenStudy (anonymous):

Thank you